To the nearest tenth, what is the area of a circle with circumference of 35.6 centimeters?
Answer by mananth(16946)

circumference = 35.6
2*pi*r=35.6

r= 35.6 / 2*pi
r= 5.66
Area of circle = pi*r^2
=pi*(5.66)^2
=100.6 square inches
Ananth